OpenStudy (anonymous):

Find the surface area of a sphere with a radius of 30 m. leave your answer in terms of pi @Directrix

Directrix (directrix):

First, the formula for the surface area of a sphere.

Directrix (directrix):

A = 4 * π * r²

Directrix (directrix):

A = 4 * π * r² A = 4 * π * 30²

OpenStudy (anonymous):

4*3.14*30^2?

Directrix (directrix):

Did you get 3600 π

OpenStudy (anonymous):

no i think i did it wrong

Directrix (directrix):

Is 3600 π among the answer options

OpenStudy (anonymous):

yes

Directrix (directrix):

Post the next question, please.